40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 20 09:32:19 2015, Steve, Update, PEM, dusty Monday morning dustyMorning.png
    Reply  Fri Feb 6 10:03:23 2015, Steve, Update, PEM, dusty days duststorm.png
       Reply  Mon Apr 27 09:09:15 2015, Steve, Update, PEM, air cond filters checked 400d.png
          Reply  Tue May 29 16:33:04 2018, Steve, Update, PEM, air cond filters replaced 
             Reply  Wed Jun 20 09:16:56 2018, Steve, Update, PEM, dusty lab AP.JPGITMY.JPGITMX.JPGSP.JPG
Message ID: 13895     Entry time: Tue May 29 16:33:04 2018     In reply to: 11259     Reply to this: 13990
Author: Steve 
Type: Update 
Category: PEM 
Subject: air cond filters replaced 

Chris replaced some air condition filters and ordered some replacement filter today.






Yesterday morning was dusty. I wonder why?

The PRM sus damping was restored this morning.

Yesterday afternoon at 4 the dust count peaked 70,000 counts

Manasa's alergy was bad at the X-end yesterday. What is going on?

There was no wind and CES neighbors did not do anything.

Air cond filters checked by Chris. The 400 days plot show 3 bad peaks at 1-20, 2-5 & 2-19


ELOG V3.1.3-